# Scalable Quantum Simulations of Scattering in Scalar Field Theory on 120 Qubits

### v1

@@ IV.4 Summary of variational methods and error budget
  • The vacuum $|\psi_{\text{vac}}\rangle$ is created using an input state preparation circuit and one step of SC-ADAPT-VQE, shown in Figs. 3a and b. The circuits are trained on $L=2-12$, and parameters are extrapolated to $L=60$. This step of the state preparation has two-qubit gate depth 25 and three variational parameters.
@@
  • The two-wavepacket state $|\psi_{\text{2wp}}\rangle$ is initialized using two layered-brickwall circuits in parallel (one layer is shown in Fig. 3c). One set of parameters is used for wavepackets with opposite momenta by spatially flipping the circuit and introducing site-wise SWAP gates. These circuits are determined using the $L=60$ vacuum parameters on an $L=12$ system, and the angles are not extrapolated. This step of the state preparation has a two-qubit gate depth of 14 and 80 parameters.
@@
  • Time evolution for $t=1-9$ is implemented using the translationally-invariant brickwall circuit shown in Fig. 6. It has a second-order Trotter step-like structure, and the number of steps is increased with time. The circuits are optimized using time evolution of the two-wavepacket state $|\psi_{\text{2wp}}\rangle$. They are trained on a lattice of size $L=12$ with the vacuum preparation parameters for $L=60$ and the wavepacket preparation parameters determined in the previous step. The parameters are not extrapolated. This part of the circuit has a two-qubit gate depth of $22\lceil\frac{t}{3}\rceil$ and 72 variational parameters.
@@
The systematic errors introduced by the approximations used in this work can be quantified by running exact classical simulations using qiskit [151]. The local infidelity $I_{d}$ (Eq. (19)), which is observed to be independent of system size as a result of the SVC framework, is used to quantify the local quality of the prepared state compared to the state of interest. The width of the region, $d$, over which the reduced state infidelity is computed, is chosen to be representative of the given step of the algorithm. For vacuum preparation, a subsystem spanning lattice sites covering several correlation lengths is used. Because of translational invariance of the vacuum, this region can be located anywhere on the lattice, and $I_{d}$ is observed to be uniform throughout. [^17] Translational invariance is seen over pairs of sites. This is a result of the vacuum preparation scheme using only operators spanning two neighboring sites. See Sec. IV.1 for a discussion. For wavepacket preparation, the wavepacket width is used. A subsystem spanning the area where the interaction takes place from $t=1-9$ is used for all steps of time evolution. The values of $I_{d}$ at each step of the simulation algorithm are shown in Table 1.
$I_{d}$ of parameters Parameter values
@@
Time evolution $(d=10)$ $t=1$ 0.1680 0.1275 72 Table 7
@@
Table 1: Values of the local infidelity $I_{d}$ for all steps of the simulation algorithm for both the free and interacting theories. For vacuum preparation, a range of 4 lattice sites is used, which is roughly 2-3×3$\times$3 × the correlation length of the state. This number is uniform throughout the lattice because of translational invariance. For wavepacket preparation, the reduced state of the wavepacket (occupying three spatial sites) is used. For time evolution, the reduced state on 10 spatial sites is used, which contains the region where the interaction takes place. The number of variational parameters used for each step is shown in the third column. The table where the corresponding parameters are specified is given in the rightmost column.
It can be seen that the infidelities for $\lambda=0$ are consistently larger than those for $\lambda=2$. This is a result of the interacting theory having a smaller correlation length than the free theory. Because circuits with equal numbers of parameters and equal two-qubit gate depths are used to initialize and time-evolve states for both values of the coupling $\lambda$, this difference in quality is expected. This is an indication that the circuit depth needs to be increased for states with larger $\xi$ in order to spread correlations over a greater number of lattice sites.
It should be noted that $I_{d}$, as well as other measures on the wavefunction, is a conservative estimate of the quality of observables calculated from these simulations. In fact, states with a modest overlap with the target state are seen to reproduce observables in the target state quite well. Appendix A shows the similarity of observables computed using exact and variational methods.
@@ V Results
$t$ of variational steps (per $t$) # of two-qubit gates Two-qubit gate depth # of PTs/layouts (per circuit) # of TREX twirls (per circuit) # of shots (per circuit)
@@
Table 2: Details of the quantum simulations performed on 120 qubits of IBM’s ibm_fez superconducting quantum computer. For a given simulation time $t$ (first column), the number of variational steps is given by $\lceil\frac{t}{3}\rceil$ (second column). The total number of two-qubit gates and corresponding two-qubit gate depth are given in columns three and four respectively, accounting for gate cancellations from different parts of the circuits and possible optimizations. Each circuit is executed using a number of Pauli-twirled circuits, with each one being mapped onto randomly-chosen rotation of the selected qubit layout; this number is given in column five. Measurement mitigation is applied using two TREX twirls for each Pauli-twirled circuit (sixth column). The number of shots per circuit is given in the rightmost column.
@@
Using the SVC framework, the circuit shown in Fig. 8 is scaled up to $L=60$ (120 qubits) and executed on ibm_fez to simulate the scattering of two wavepackets in the $\phi^{4}$ theory. $n_{q}=2$ qubits and a cutoff of $\phi_{\text{max}}=1.5$ are used to represent the state of the field at each spatial site with a bare mass of $m=1/2$. Error mitigation is vital for the successful extraction of observables from noisy quantum simulations [9]. An overview of the error mitigation methods used in this work is given here, and the implementation details and choices can be found in App. C. Once the bare circuits are created, several error mitigation layers are added. Dynamical decoupling [152, 153] is used to mitigate idle errors and crosstalk between qubits. Pauli twirling (PT) [154] is added to all two-qubit gates (CZ for ibm_fez) to convert coherent errors to incoherent stochastic noise, after which a depolarizing noise model is assumed for each observable of interest. Similar to PT for two-qubit gates, Twirled Readout Error eXtinction (TREX) [155] is used to mitigate measurement errors. Observables are then estimated using Operator Decoherence Renormalization (ODR) [10, 15, 156, 157], which involves running a “mitigation” circuit for every Pauli- and TREX-twirled “physics” circuit. Circuits used for mitigation must be classically simulable, but have similar error profiles to the physics circuits. In this work, circuits implementing the time evolution of the vacuum state $e^{-itH}|\psi_{\text{vac}}\rangle$ are used as mitigation circuits. Although the vacuum preparation and time evolution are implemented approximately and $e^{-itH}|\psi_{\text{vac}}\rangle\neq|\psi_{\text{vac}}\rangle$, the vacuum time evolution can be determined classically via an exponential extrapolation. See App. C for details on the classical calculation of this time evolution. Compared to methods that have been used in the past, it is found that the vacuum evolution circuits more accurately reflect the noise in the wavepacket circuits. This is especially noticeable for the case of brickwall circuits, where the gates are densely packed and noise spreads faster than in Trotter time evolution circuits. Using the mitigation circuits, the depolarization parameters are computed using the ratio of the measured outcome $\langle O_{j}\rangle_{\text{meas}}$ to the expected noiseless outcome $\langle O_{j}\rangle_{\text{true}}$ for each local observable $O_{j}$
@@
[…]
@@
Figure 10: A detailed view of the vacuum-subtracted $\langle\phi^{2}_{j}\rangle$ for the scattering of two wavepackets, as shown in Fig. 9, by time step. Noiseless classical MPS simulations (grey dashed lines) are compared to data from IBM’s quantum computer ibm_fez after error mitigation (points) for $\lambda=0$ (left column) and $\lambda=2$ (right column). In the device data, the vacuum evolution is evaluated via a classically-determined exponential extrapolation. The error bars represent one standard deviation of the bootstrap-resampled data.
@@
The two-qubit gate depths and counts, as well as other parameters of the runs on ibm_fez for times $t=1-9$, are shown in Table 2. Circuits with a maximum two-qubit gate depth of 103 using 4924 two-qubit gates (for $t=7,8,9$) are used. For time $t$, $\lceil\frac{t}{3}\rceil$ variational steps are used (the variational step is defined in Sec. IV.3 and shown in Fig. 6). For each time $t$, four circuits are executed: $e^{-itH}|\psi_{\text{2wp}}\rangle$ and $e^{-itH}|\psi_{\text{vac}}\rangle$ for $\lambda=0$ and $\lambda=2$. 80 Pauli-twirled circuits are run for each physics and mitigation circuit. For each Pauli-twirled circuit, two TREX twirled circuits are run for measurement mitigation. Each circuit executed on the device is evaluated using 8000 shots.
@@
In these simulations, fluctuations of $\langle\phi^{2}_{j}\rangle$ above the vacuum are measured on the quantum device. [^18] Note that it is possible to efficiently measure $\langle H_{j}\rangle$ using the same setup. Besides $\Pi^{2}_{j}$, all terms in the Hamiltonian of Eq. (1) are in the $\phi$ basis, so their expectation values can be computed using measurements in the $\phi$ basis. To measure in the $\Pi$ basis, a second set of circuits can be run with the $k_{\phi}$ local Fourier transform appended. The results can then be added together to compute $\langle H_{j}\rangle$. From this, it is straightforward to estimate quantities that are accessible to analytic computations, such as cross sections computed using perturbative quantum field theory. The results from the device and from MPS simulations for $\langle\phi^{2}_{j}\rangle_{\text{2wp}}-\langle\phi^{2}_{j}\rangle_{\text{vac}}$ for $\lambda=0$ and $\lambda=2$ are shown in Figs. 9(a) and 9(b) respectively. A breakdown of the results by time step is shown in Fig. 10, and numerical values for these runs are given in Tables 16-24. The uncertainties in the results from the quantum computer are estimated using bootstrap resampling. The expected results are determined using the qiskit [151] MPS simulator with a maximum bond dimension of 100. For the circuits considered, it is seen that the results of these simulations are converged to $10^{-2}-10^{-3}$.
@@
The scattering process is seen by examining $\langle\phi^{2}_{j}\rangle_{\text{2wp}}-\langle\phi^{2}_{j}\rangle_{\text{vac}}$ in Fig. 9. The propagation of particles is clearly identified as disturbances in the vacuum-subtracted $\langle\phi^{2}_{j}\rangle$ in the interaction region in the center (spatial sites 24-36). These are indicated as bright spots on the plots. The wavepackets, centered to the left and right of the center of the lattice at $t=0$, are initialized with opposite momenta. They travel toward each other and collide in the center. After the collision takes place, a light cone is seen to develop, indicating that the particles have propagated past each other and are travelling away from the point of the collision.
@@
The effect of interactions is clearly seen when comparing results in the interaction region between the free and interacting theories as in Fig 11. In the free theory, the collision is seen to peak around $t=4$, after which the particles travel away from each other. When interactions are included in the theory, the peak of the collision is observed later (around $t=5$), and a time delay due to the interaction is visible. The difference between $\lambda=0$ and $\lambda=2$ can also be seen at late times $(t=9)$, where the free particles can be seen having travelled one more spatial site than the interacting particles. Because the $\phi^{4}$ potential is repulsive, particles with overlapping wavefunctions incur an energy penalty. In this sense, the time delay can be viewed as the interacting particles encountering a potential step and thus losing kinetic energy when their wavepackets overlap. As a result, the particles spend a longer time in the central region where they are able to interact. It is expected that this effect will be even more prominent at greater interaction strengths.
@@
Figure 11: A zoomed-in view of the interaction region of Fig. 9 highlighting the effect of interactions. The vacuum-subtracted $\langle\phi_{j}^{2}\rangle$ in the free theory (left panels) is compared to the interacting theory (right panels). The top panels show error-mitigated results from IBM’s quantum computer ibm_fez, while the bottom panels show classical noiseless MPS simulations.
@@
The results in Figs. 9, 10, and 11 show that by using a combination of variational circuit methods (SC-ADAPT-VQE and brickwall), together with error mitigation techniques, quantum computers are able to simulate scattering processes of interacting particles. The scalability of each step of this simulation algorithm enables the simulation of collision events on arbitrarily large system sizes in a systematically-improvable manner. It should be noted that the observation of $\langle\phi^{2}_{j}\rangle_{\text{2wp}}-\langle\phi^{2}_{j}\rangle_{\text{vac}}=0$ in the vacuum regions is a nontrivial result, as it requires the wavepacket simulations from the quantum computer to agree closely with extrapolations from classical computations. In Figs. 9 and 10, the data from the quantum computer is seen to qualitatively agree with the MPS simulations. Increasing levels of noise (as well as uncertainties) can be seen in the results for later times, which is expected as a result of deeper circuits. There exist points where the error-mitigated results still disagree from the MPS data by several standard deviations (see Fig. 10). This is particularly noticeable in the interaction region during the collision of the particles, but occurs in vacuum regions as well. This highlights a weakness of the chosen error mitigation method: because the state chosen for error mitigation is $e^{-itH}|\psi_{\text{vac}}\rangle$, its evolution does not capture potential changes to wavepackets. Because the noise is state-dependent, these mitigation circuits cannot mimic the noise in the circuits simulating scattering perfectly. This reflects the simplicity of the assumed error model, and is an area where much improvement is expected in the future. Although errors persist even after the application of error mitigation, it is clear that quantum computers today are capable of probing scattering events, potentially shedding light on inelastic collisions in more complicated theories in the future.

### v2

@@ IV.4 Summary of variational methods and error budget
  • The vacuum $|\psi_{\text{vac}}\rangle$ is created using an input state preparation circuit and one step of SC-ADAPT-VQE, shown in Figs. 3a and b. The circuits are trained on $|\psi_{\text{targ}}\rangle$ determined from ED with L=212L=2-12, and parameters are extrapolated to $L=60$. This step of the state preparation has a two-qubit gate depth of 25 and three variational parameters.
@@
  • The two-wavepacket state $|\psi_{\text{2wp}}\rangle$ is initialized using two layered brickwall circuits in parallel (one layer is shown in Fig. 3c). A single set of parameters is used for wavepackets with opposite momenta by spatially flipping the circuit and introducing site-wise SWAP gates. These circuits are determined using the $L=60$ vacuum parameters on an $L=12$ system, and the angles are not extrapolated. The prepared wavepackets span three lattice sites and have $k=\pm\pi/3$ and $\sigma_{k}=\pi/3$. The continuum expression for the $\lambda=0$ wavepacket (Eq. (10)) and the resulting state after the adiabatic turn-on of $\lambda$ (Eq. (12)) are used as $|\psi_{\text{targ}}\rangle$ for the free and interacting theories, respectively. This step of the state preparation has a two-qubit gate depth of 14 and 80 parameters.
@@
  • Time evolution for t=19t=1-9 is implemented using the translationally-invariant brickwall circuit shown in Fig. 6. It has a second-order Trotter step-like structure, and the number of steps is increased with time. The circuits are optimized using time evolution of the two-wavepacket state $|\psi_{\text{2wp}}\rangle$ as $|\psi_{\text{targ}}\rangle$, determined using 100 second-order Trotter steps. They are trained on a lattice of size $L=12$ with the vacuum preparation parameters for $L=60$ and the wavepacket preparation parameters determined in the previous step. The parameters are not extrapolated. This part of the circuit has a two-qubit gate depth of $22\lceil\frac{t}{3}\rceil$ and 72 variational parameters.
@@
The systematic errors introduced by the variational approximations used in this work can be quantified by running exact classical simulations using qiskit [158] (see App. A for details on the digitization effects). The local infidelity $I_{d}$ (Eq. (19)), which is observed to be independent of system size as a result of the SVC framework, is used to quantify the local quality of the prepared state compared to the state of interest. The width of the region, $d$, over which the reduced state infidelity is computed, is chosen to be representative of the given step of the algorithm. For vacuum preparation, a subsystem spanning lattice sites covering several correlation lengths is used. Because of the translational invariance of the vacuum, this region can be located anywhere on the lattice, and $I_{d}$ is observed to be uniform throughout. [^19] Translational invariance is seen over pairs of sites. This is a result of the vacuum preparation scheme using only operators spanning two neighboring sites. See Sec. IV.1 for a discussion. For wavepacket preparation, the wavepacket width is used. A subsystem spanning the area where the interaction takes place from t=19t=1-9 is used for all steps of time evolution. The values of $I_{d}$ at each step of the simulation algorithm are shown in Table 1.
$I_{d}$ # of parameters Parameter values
@@
Time
evolution
$(d=10)$
$t=1$ 0.1680 0.1275 72 Table 8
@@
Table 1: Values of the local infidelity $I_{d}$ for all steps of the simulation algorithm for both the free and interacting theories. The infidelity is computed using the target states $|\psi_{\text{targ}}\rangle$ specified at the beginning of this section. For vacuum preparation, a range of 4 lattice sites is used, which is roughly 2-$3\times$ the correlation length of the state. The local infidelity is uniform throughout the lattice because of translational invariance. For wavepacket preparation, the reduced state of the wavepacket (occupying three spatial sites) is used. For time evolution, the reduced state on 10 spatial sites is used, which contains the region where the interaction takes place. The number of variational parameters used for each step is shown in the third column. The table where the corresponding parameters are specified is given in the rightmost column.
It can be seen that the infidelities for $\lambda=0$ are consistently larger than those for $\lambda=2$. This is a result of the interacting theory having a smaller correlation length than the free theory. Because circuits with equal numbers of parameters and equal two-qubit gate depths are used to initialize and time-evolve states for both values of $\lambda$, this difference in quality is expected. This is an indication that the circuit depth needs to be increased for states with larger $\xi$ in order to spread correlations over a greater number of lattice sites.
It should be noted that $I_{d}$, as well as other measures on the wavefunction, is a conservative estimate of the quality of observables calculated from these simulations. In fact, states with a modest overlap with the target state are seen to reproduce observables in the target state quite well. Appendix C shows the similarity of observables computed using exact and variational methods.
@@ V Results
$t$
# of variational steps (per $t$) # of two-qubit gates Two-qubit gate depth # of PTs/layouts (per circuit) # of TREX twirls (per circuit) # of shots (per circuit)
@@
Table 2: Details of the quantum simulations performed on 120 qubits of IBM’s ibm_fez superconducting quantum computer. For a given simulation time $t$ (first column), the number of variational steps is given by $\lceil\frac{t}{3}\rceil$ (second column). The total number of two-qubit gates and corresponding two-qubit gate depths are given in columns three and four respectively, accounting for gate cancellations from different parts of the circuits and possible optimizations. Each circuit is executed using a number of Pauli-twirled circuits, with each one being mapped onto randomly-chosen rotation of the selected qubit layout; this number is given in column five. Measurement mitigation is applied using two TREX twirls for each Pauli-twirled circuit (sixth column). The number of shots per circuit is given in the rightmost column.
@@
Using the SVC framework, the circuit shown in Fig. 8 is scaled up to $L=60$ (120 qubits) and executed on ibm_fez to simulate the scattering of two wavepackets in the $\phi^{4}$ theory. The theory is digitized onto $n_{q}=2$ qubits with a cutoff of $\phi_{\text{max}}=1.5$ to represent the state of the field at each spatial site with a bare mass of $m=1/2$. Error mitigation is vital for the successful extraction of observables from noisy quantum simulations [9]. An overview of the error mitigation methods used in this work is given here, and the implementation details and choices can be found in App. E. Once the bare circuits are created, several error mitigation layers are added. Dynamical decoupling (DD) [159, 160] is used to mitigate idle errors and crosstalk between qubits. Pauli twirling (PT) [161] is added to all two-qubit gates (CZ for ibm_fez) to convert coherent errors to incoherent stochastic noise, after which a depolarizing noise model is assumed for each observable of interest. Similar to PT for two-qubit gates, Twirled Readout Error eXtinction (TREX) [162] is used to mitigate measurement errors. Observables are then estimated using Operator Decoherence Renormalization (ODR) [10, 15, 163, 164], which involves running a “mitigation” circuit for every Pauli- and TREX-twirled “physics” circuit. Circuits used for mitigation must be classically simulable, but must have similar error profiles to the physics circuits. In this work, circuits implementing the time evolution of the vacuum state $e^{-itH}|\psi_{\text{vac}}\rangle$ are used as mitigation circuits. Although the vacuum preparation and time evolution are implemented approximately and $e^{-itH}|\psi_{\text{vac}}\rangle\neq|\psi_{\text{vac}}\rangle$, the vacuum time evolution can be determined classically via an exponential extrapolation. See App. E for details on the classical calculation of this time evolution. Compared to methods that have been used in the past, it is found that the vacuum evolution circuits more accurately reflect the noise in the wavepacket circuits. This is especially noticeable for the case of brickwall circuits, where the gates are densely packed and noise spreads faster than in Trotter time evolution circuits. Using the mitigation circuits, the depolarization parameters are computed using the ratio of the measured outcome $\langle O_{j}\rangle_{\text{meas}}$ to the expected noiseless outcome $\langle O_{j}\rangle_{\text{true}}$ for each local observable $O_{j}$
@@
Figure 9: The time evolution of the vacuum-subtracted $\langle\phi^{2}_{j}\rangle$ for a $L=60$ (120 qubits) lattice for (a) the free $(\lambda=0)$ and (b) interacting $(\lambda=2)$ theories. Two wavepackets are initialized with opposite momenta $(k=\pm\pi/3)$ at $t=0$ and centered at sites j=28, 32j=28,\,32. The particles travel toward each other, collide, and propagate as time goes on. Error-mitigated results from IBM’s quantum computer ibm_fez (top panels) are compared to noiseless MPS circuit simulations (bottom panels). The vacuum of the theory is built using variational circuits determined by SC-ADAPT-VQE, while the wavepacket preparation and time evolution is implemented by variational brickwall circuits. In the device data, the vacuum evolution is evaluated via a classically-determined exponential extrapolation. A broken-down view for each time slice is shown in Fig. 10, and a zoomed-in comparison of the scattering region between the free and interacting theories is given in Fig. 11. Details on the error mitigation strategies are given in the main text and in App. E.
@@
Figure 10: A detailed view of the vacuum-subtracted $\langle\phi^{2}_{j}\rangle$ for the scattering of two wavepackets, as shown in Fig. 9, by time step. Noiseless MPS circuit simulations (grey dashed lines) are compared to data from IBM’s quantum computer ibm_fez after error mitigation (points) for $\lambda=0$ (left column) and $\lambda=2$ (right column). In the device data, the vacuum evolution is evaluated via a classically-determined exponential extrapolation. The error bars represent one standard deviation of the bootstrap-resampled data.
@@
The two-qubit gate depths and counts, as well as other parameters of the runs on ibm_fez for times t=19t=1-9, are shown in Table 2. Circuits with a maximum two-qubit gate depth of 103 using 4924 two-qubit gates (for $t=7,8,9$) are used. For time $t$, $\lceil\frac{t}{3}\rceil$ variational steps are used (the variational step is defined in Sec. IV.3 and shown in Fig. 6). For each time $t$, four circuits are executed: $e^{-itH}|\psi_{\text{2wp}}\rangle$ and $e^{-itH}|\psi_{\text{vac}}\rangle$ for $\lambda=0$ and $\lambda=2$. 80 Pauli-twirled circuits are run for each physics and mitigation circuit. For each Pauli-twirled circuit, two TREX twirled circuits are run for measurement mitigation. Each circuit executed on the device is evaluated using 8000 shots.
@@
In these simulations, fluctuations of $\langle\phi^{2}_{j}\rangle$ above the vacuum are measured on the quantum device. [^20] Note that it is possible to efficiently measure $\langle H_{j}\rangle$ using the same setup. Besides $\Pi^{2}_{j}$, all terms in the Hamiltonian of Eq. (1) are in the $\phi$ basis, so their expectation values can be computed using measurements in the $\phi$ basis. To measure in the $\Pi$ basis, a second set of circuits can be run with the $k_{\phi}$ local Fourier transform appended. The results can then be added together to compute $\langle H_{j}\rangle$. From this, it is straightforward to estimate quantities that are accessible to analytic computations, such as cross sections computed using perturbative quantum field theory. The results from the device and from MPS circuit simulations for $\langle\phi^{2}_{j}\rangle_{\text{2wp}}-\langle\phi^{2}_{j}\rangle_{\text{vac}}$ for $\lambda=0$ and $\lambda=2$ are shown in Figs. 9(a) and 9(b) respectively. A breakdown of the results by time step is shown in Fig. 10, and numerical values for these runs are given in Tables 17-25. The uncertainties in the results from the quantum computer are estimated using bootstrap resampling. The expected results are determined using the qiskit [158] MPS circuit simulator. This classical method works by converting the circuit to a tensor network acting on a MPS and applying the circuit to the MPS gate by gate while enforcing the cutoff and maximum bond dimension constraints at each step. See Ref. [165] for more information. In this work, a maximum bond dimension of 100 is used. For the circuits considered, it is seen that the results of these simulations are converged to $10^{-2}-10^{-3}$.
@@
The scattering process is seen by examining $\langle\phi^{2}_{j}\rangle_{\text{2wp}}-\langle\phi^{2}_{j}\rangle_{\text{vac}}$ in Fig. 9. The propagation of particles is clearly identified as disturbances in the vacuum-subtracted $\langle\phi^{2}_{j}\rangle$ in the interaction region in the center (spatial sites 24-36). These are indicated as bright spots on the plots. The wavepackets, centered to the left and right of the center of the lattice at $t=0$, are initialized with opposite momenta. They travel toward each other and collide in the center. After the collision takes place, a light cone is seen to develop, indicating that the particles have propagated past each other and are traveling away from the point of the collision.
@@
The effect of interactions is clearly seen when comparing results in the interaction region between the free and interacting theories as in Fig. 11. In the free theory, the collision is seen to peak around $t=4$, after which the particles travel away from each other. When interactions are included in the theory, the peak of the collision is observed later (around $t=5$), and a time delay due to the interaction is visible. The difference between $\lambda=0$ and $\lambda=2$ can also be seen at late times $(t=9)$, where the free particles can be seen having traveled one more spatial site than the interacting particles. A time delay is expected because the $\phi^{4}$ potential is repulsive, particles with overlapping wavefunctions incur an energy penalty. In this sense, the time delay can be viewed as the interacting particles encountering a potential step and thus losing kinetic energy when their wavepackets overlap. As a result, the particles spend a longer time in the central region where they are able to interact. It is expected that this effect will be even more prominent at greater interaction strengths. In the case of the digitized simulations done in this work with $n_{q}=2$, the interactions and time delay are caused by digitization effects. See App. A for a detailed examination of these interactions.
@@
Figure 11: A zoomed-in view of the interaction region of Fig. 9 highlighting the effect of interactions. The vacuum-subtracted $\langle\phi_{j}^{2}\rangle$ in the free theory (left panels) is compared to the interacting theory (right panels). The top panels show error-mitigated results from IBM’s quantum computer ibm_fez, while the bottom panels show noiseless MPS circuit simulations.
@@
The results in Figs. 9, 10, and 11 show that by using a combination of variational circuit methods (SC-ADAPT-VQE and brickwall), together with error mitigation techniques, quantum computers are able to simulate scattering processes of interacting particles. The scalability of each step of this simulation algorithm enables the simulation of collision events on arbitrarily-large system sizes in a systematically-improvable manner. It should be noted that the observation of $\langle\phi^{2}_{j}\rangle_{\text{2wp}}-\langle\phi^{2}_{j}\rangle_{\text{vac}}=0$ in the vacuum regions is a nontrivial result, as it requires the wavepacket simulations from the quantum computer to agree closely with extrapolations from classical computations. In Figs. 9 and 10, the data from the quantum computer is seen to qualitatively agree with the MPS circuit simulations. Increasing levels of noise (as well as uncertainties) can be seen in the results for later times, which is expected as a result of deeper circuits. There exist points where the error-mitigated results still disagree from the MPS data by several standard deviations (see Fig. 10). This is particularly noticeable in the interaction region during the collision of the particles, but occurs in vacuum regions as well. This highlights a weakness of the chosen error mitigation method: because the state chosen for error mitigation is $e^{-itH}|\psi_{\text{vac}}\rangle$, its evolution does not capture potential changes to wavepackets. Since the noise is state-dependent, these mitigation circuits cannot mimic the noise in the circuits simulating scattering perfectly. This reflects the simplicity of the assumed error model, and is an area where much improvement is expected in the future. Although errors persist even after the application of error mitigation, it is clear that quantum computers today are capable of probing scattering events, potentially shedding light on inelastic collisions in more complicated theories in the future.